PHASE SHIFT CAVITY RING DOWN (PS-CRD) SPECTROSCOPY AT CRYOGENIC TEMPERATURES.

\underline{ERNEST K. LEWIS}, CARLOS MANZANARES I., Department of Chemistry and Biochemistry, Baylor University, Waco, TX 76798; and MARY C. SALAZAR, ANTONIO J. HERNANDEZ, Department of Chemistry, Simon Bolivar University, Caracas 1080A, Venezuela.

Phase Shift Cavity Ring Down (PS-CRD) spectroscopy was used for the detection of high vibrational overtones in a low temperature environment. The PS-CRD technique at low temperature involves an optical cavity (static cell) attached to a cryostat. The cell and mirrors of the optical cavity are at the same temperature. A continuous wave titanium sapphire laser system in the wavelength range from 700 to 820nm and with a 0.1 cm^{−1} resolution was electro-optically modulated, and passed through the optical cavity. The change in the phase angle of the modulated laser beam after passing through the cavity was recorded as a function of the wavelength. The tangent of the phase shift angle is related to the light relaxation time. The rotationally resolved C-H overtone (v = 0 → 5) of CHD_{3} was then studied over the temperature range 296K to 100K.